OpenStudy (df001):

For each value of Θ, use half angle formula find Sin(1/2)Θ, Cos(1/2)Θ, tan (1/2)Θ

OpenStudy (df001):

\[\cos \frac{ 1 }{ 2 }\theta = \pm \sqrt{\frac{ 1+\theta }{ 2 }}\]

OpenStudy (df001):

help me with Sin(1/2)x please!!